A charge of 240 coulombs passes through the heater in one minute. What current flows in the heater?
1
Expert's answer
2018-08-15T09:01:08-0400
We can find the amount of the current that flows in the heater from the formula:
I=Q/t=(240 C)/(60 s)=4 A.